
Coin Toss Lab
Presentation
•
Biology
•
12th Grade
•
Practice Problem
•
Hard
Alicia Archer
FREE Resource
10 Slides • 1 Question
1
Open Ended
If you flip a coin, what is the chance it will come up heads?
2
You can either get heads or tails. You will
get one of two results. Therefore, the
probability of getting heads is one out of
two. One out of two is 1/2, one half, or 50%!
3
Flipping 2 Coins
If you flipped two coins at the same time, what is the probability that both coins will be heads? Probability is the chance or likelihood something is going to happen. There is a random chance to be heads or tails, meaning it is equally likely to get either outcome.
4
HH
HT
TH
TT
5
There are four possible outcomes for flipping 2 coins: HH, HT, TH, and TT.
Only one outcome has both heads out of four. The chance of flipping 2 coins and getting both heads is therefore one out of four. 1/4, one fourth, or 25%. To calculate percent you divide your fracton, one divided by four, and multiple by 100%.
1/4 = .25
.25 x 100% = 25%
6
Another way to calculate this is multiplying 1/2 by 1/2 or multiplying .5 by .5. You can also take your chance and divide it by 2 on your calculator for every coin flip. I will demonstrate using the calculator on my computer.
7
Please click the following link to make a copy of the assignment:
Change the title to include your last name!
8
Change Your Share Settings
Click Share on the top right > go to general settings > change your share settings to "Clark County School District". If you don't I won't be able to see your assignment when you turn it in!
9
The Rules are:
No coin touches the floor
No coin travels through the air towards another student.
The tosses that you should make and record are:
One coin for 50 tosses. Record both heads and tails. It helps if one person counts the number of tosses and the second person records the results of each toss.
Two coins tossed at the same time for 25 trials. Only record when both coins land heads up! Ignore all other results.
Three coins tossed at the same time for 25 trials. Only record when all three coins land heads up! Ignore all other results.
10
Calculator
You can use the calculator on your chomebook. You can use your phone calculator ONLY if you use it as a calculator. If I see you using it for anything else I will ask you to put it away!
11
How to Graph Using Sheets
Watch me create a graph using Google sheets so you can insert a graph into your assignment.
If you flip a coin, what is the chance it will come up heads?
Show answer
Auto Play
Slide 1 / 11
OPEN ENDED
Similar Resources on Wayground
10 questions
MATERI 2 - MEETING 3 KELAS XII
Presentation
•
12th Grade
6 questions
Sesión 5 Razonamiento Verbal
Presentation
•
12th Grade
10 questions
Unit 1 Review
Presentation
•
12th Grade
8 questions
Quantum Mechanics
Presentation
•
University
10 questions
PROBABILITY
Presentation
•
University
10 questions
Koala
Presentation
•
University
8 questions
Salvadoran Ethnic Night
Presentation
•
12th Grade
12 questions
Herança Ligada ao Sexo - 3º ano
Presentation
•
12th Grade
Popular Resources on Wayground
10 questions
Safety Day 2026 – Student Review
Quiz
•
9th Grade
21 questions
Continents and Oceans
Quiz
•
6th Grade
20 questions
Parts of Speech
Quiz
•
5th Grade
16 questions
Subject & Predicate
Quiz
•
5th Grade
25 questions
Multiplication Facts
Quiz
•
5th Grade
12 questions
Map Skills
Quiz
•
3rd Grade
22 questions
Continents and Oceans
Quiz
•
5th Grade
15 questions
Rounding Decimals
Quiz
•
5th Grade
Discover more resources for Biology
20 questions
Cell Transport
Quiz
•
9th - 12th Grade
8 questions
DNA vs. RNA
Presentation
•
9th - 12th Grade
16 questions
Epithelial tissues
Quiz
•
11th - 12th Grade
15 questions
DNA Structure and Replication
Quiz
•
9th - 12th Grade
20 questions
Ecological Succession Review
Quiz
•
9th - 12th Grade
3 questions
Enzymes & Energy
Quiz
•
9th - 12th Grade
38 questions
Mitosis & Meiosis
Quiz
•
9th - 12th Grade
37 questions
DNA/RNA practice part 2
Quiz
•
9th - 12th Grade